# Nonparametric Estimation of Matching Efficiency and Elasticity in a Marriage Agency Platform: 2014–2025

### v1

@@ Abstract
This paper examines monthly matching efficiency in the Japanese marriage market using novel data from IBJ, the country’s largest structured matching platform. Unlike administrative or dating app data, IBJ provides full search, dating, and matching logs based on verified profiles and confirmed engagements. Users are highly selected into serious marriage search via costly screening. Covering 3.3% of national marriages annually, the data offer rare behavioral granularity. Using a nonparametric approach, I estimate time-varying matching functions and elasticities. Efficiency rises fourfold over time, illustrating how digital intermediation transforms partner search in modern marriage markets.
@@ 1 Introduction
Understanding how individuals find partners in modern marriage markets is central to the study of family economics (Chiappori and Salanié 2023). As marriage rates decline and matching technologies proliferate, questions surrounding search efficiency, partner availability, and market design have grown increasingly salient. While theoretical models of two-sided search abound, empirical research has been hampered by data limitations: traditional sources such as administrative records or surveys capture only realized marriages, omitting the entire search and matching process that precedes them, as well as the presence of singles who are not seriously seeking partners. Consequently, little is known about the dynamics of partner search, how modern platforms shape matching outcomes, or how efficiently modern markets operate. This study addresses these limitations by using novel data from IBJ, Japan’s largest structured marriage matching platform, which accounts for approximately 3.3% of all marriages in Japan in 2024. The dataset includes over 10,000 confirmed engagements annually and allows for behavioral analysis of the full matching process, including search, proposals, and formal engagements. In this note, we focus on documenting macro-level trends on the platform as a first step toward more detailed analysis, leaving in-depth investigations of user behavior and underlying mechanisms to future work.
@@
Using this dataset, I estimate a matching function following the nonparametric approach of Lange and Papageorgiou (2020) to evaluate the platform’s performance. [^1] Their approach is widely applied to labor market settings. See Otani (2024a, b) and Kanayama and Otani (2024). I find that matching efficiency increased fourfold between 2014 and 2025, with sharp gains beginning in 2017, coinciding with major platform consolidation. Engagement formation also became more responsive to user composition, as indicated by rising elasticities. These results underscore the role of digital intermediation in shaping modern family formation. Unlike traditional arrangements mediated by kinship networks or local norms, platforms like IBJ provide structured, data-rich environments where frictions are minimized, outcomes are verified, and behavioral dynamics are observable, enabling new empirical insights.
@@ 2 Data
I use confidential data from IBJ covering the period 2014–2025. In 2024, IBJ accounted for 3.3% of all marriages in Japan, with over 10,000 engagements annually. While this share is smaller than administrative marriage registers, IBJ offers unique visibility into the pre-marriage process. The platform collects verified user data, including age, income, education, and marital status, and tracks the full sequence of user actions: search, proposals, messages, and dates. All records are time-stamped and linked to observed engagement outcomes. IBJ users are highly selected into serious long-term partnership search. Entry requires screening and significant upfront, monthly, and exit fees, which effectively exclude casual users. In contrast to dating apps that rely on self-reported profiles and open-ended interactions, IBJ functions as a closed, consultant-mediated two-sided matching platform where mutual consent leads to a confirmed engagement. This structured environment makes the data well-suited for analyzing matching efficiency and responsiveness over time.
@@
Panels (a) and (b) in Figure 1 display trends in the number of active users and engagements on the IBJ platform from 2014 to 2025. Both female $F$ and male $M$ participation rise steadily, with male user growth outpacing that of females—particularly after 2020—leading to a gradual decline in market tightness, measured as the gender balance ratio $\frac{F}{M}$. Alongside this compositional shift, the number of engagements $E$ remains relatively flat until around 2020, after which it increases sharply. This acceleration likely reflects improvements in platform performance, such as better algorithmic matching, increased user activity, or broader social adoption of digital matching services.
Panel (c) illustrates the partner finding rates for female and male users per month, measured respectively as $\frac{E}{F}$ and $\frac{E}{M}$. Both rates increase over time, with females consistently achieving higher matching rates. This suggests improved match formation efficiency on the platform, particularly for female users, despite the growing supply of male participants. The male rate displays greater month-to-month volatility, pointing to more sensitivity to changes in market conditions.
@@ 3 Empirical Framework
Let $E_{t}$ denote the number of engagements at time $t$, $F_{t}$ and $M_{t}$ the number of active female and male users, and $A_{t}$ a time-varying matching efficiency parameter. Let $(A,F,M)$ denote random variables corresponding with realizations subscripted by time $t$. Let assume that $M$ and $A$ are independent conditional on $F$, that is $M\mathop{\perp}A|F$. I assume a matching function of the form $E_{t}=m(A_{t}F_{t},M_{t})$ with constant returns to scale. Given the assumptions, by applying the nonparametric identification results of Matzkin (2003) to avoid functional form assumptions, Proposition 1 of Lange and Papageorgiou (2020) shows that the observed joint distribution of engagements, female users, and male users, denoted as $G(E,F,M)$, nonparametrically identifies the joint distribution of matching efficiency and female participation, $H(A,F)$, as well as the matching function $m(AF,M):\mathbb{R}^{2}_{+}\rightarrow\mathbb{R}_{+}$, up to a normalization of the efficiency parameter $A$ at a given point $A_{0}$ in the support of $(A,F,M)$.
@@ 4 Results
Using nonparametric methods proposed by Lange and Papageorgiou (2020), I estimate time-varying matching efficiency $A_{t}$ and elasticities with respect to effective female search input $A_{t}F_{t}$ and male users $M_{t}$. [^3] Details are shown in Appendix. Finite sample performance via Monte Carlo simulations is also shown in Otani (2024b).
@@ 4.1 Matching efficiency and elasticity in the marriage platform
In Figure 2, matching efficiency increases more than fourfold from baseline levels in 2014, with particularly rapid acceleration beginning around 2017 and continuing through the post-2020 period. Notably, the efficiency nearly triples between 2017 and 2019. This sharp improvement may be partly attributable to structural changes within the platform, including the consolidation of other major marriage agencies into the IBJ network. For example, the integration of ZWEI—another nationwide marriage consultation service—into the IBJ Group during this period likely expanded the effective user base and improved cross-platform coordination. Such developments may have enhanced the breadth of matching opportunities and the sophistication of recommendation mechanisms, thereby contributing to substantial gains in match formation efficiency. Elasticity with respect to female-side input ranges from 0.5 to 1.2, while male-side elasticity frequently exceeds 1.0, indicating that the platform has become increasingly responsive to shifts in both sides of the market.
@@
Together, these trends indicate that the IBJ platform has become not only more efficient in producing engagements but also more responsive to shifts in both sides of the user base. The increase in efficiency and elasticity suggests improved functionality in the underlying matching process, possibly driven by better recommendation algorithms, increased user base scale, or more targeted user engagement strategies.
4.2 Regional differences in matching efficiency and elasticity
@@ 5 Conclusion
This paper documents substantial improvements in matching efficiency on a large-scale marriage platform in Japan. Leveraging verified user data and well-defined engagement outcomes, I apply a nonparametric approach to estimate time-varying matching functions. Efficiency and elasticity estimates indicate that the platform has become more effective and responsive to user-side dynamics. Regional analysis confirms consistent functionality across major urban areas. These findings highlight the potential of digital matching platforms to reshape traditional partner search markets through structured intermediation and algorithmic improvements.
@@ A.1 Estimation details
where $\kappa(\cdot)$ is a bivariate normal kernel function with bandwidth set to 0.01.
@@
Finally, I compute local elasticities of the matching function by regressing engagements $E$ on female and male user counts, interacted with the estimated effective search input $AF$. This allows us to estimate the marginal responsiveness of engagements to changes in female and male participation:
[…]

### v3

@@ Abstract
This paper examines monthly matching efficiency in the Japanese marriage market using novel data from IBJ, the country’s largest structured matching platform. Unlike administrative or dating app data, IBJ provides full search, dating, and matching logs based on verified profiles and confirmed engagements. Users are highly selected into serious marriage search via costly screening. Covering 3.3% of national marriages annually, I estimate time-varying matching functions using a nonparametric approach. Matching efficiency triples between 2014 and 2025, with diverging elasticities indicating increasingly asymmetric responsiveness. Results show how digital intermediation reshapes partner search in modern marriage markets.
@@ 1 Introduction
Understanding how individuals find partners in modern marriage markets is central to the study of family economics (Chiappori and Salanié 2023). As marriage rates decline and matching technologies proliferate, questions surrounding search efficiency, partner availability, and market design have grown increasingly salient. While theoretical models of two-sided search abound, empirical research has been hampered by data limitations: traditional sources such as administrative records or surveys capture only realized marriages, omitting the entire search and matching process that precedes them, as well as the presence of singles who are not seriously seeking partners. Consequently, little is known about the dynamics of partner search, how modern platforms shape matching outcomes, or how efficiently modern markets operate. This study addresses these limitations by using novel data from IBJ, Japan’s largest structured marriage matching platform, which accounts for approximately 3.3% of all new marriages in Japan in 2024.111Approximately 50% of engagements among IBJ users occur within the platform, while the remaining 50% result from relationships formed outside the platform. The dataset includes over 10,000 confirmed engagements annually and allows for behavioral analysis of the full matching process, including search, proposals, and formal engagements. In this note, we focus on documenting macro-level trends on the platform as a first step toward more detailed analysis, leaving in-depth investigations of user behavior and underlying mechanisms to future work.
@@
Using this dataset, I estimate a matching function following the nonparametric approach of Lange and Papageorgiou (2020) to evaluate the platform’s performance. [^2] Their approach is widely applied to labor market settings. See Otani (2024a, b) and Kanayama and Otani (2024). I find that matching efficiency increased threefold between 2014 and 2025, with gains beginning in 2017, coinciding with major platform consolidation. Engagement formation also became more responsive—though increasingly asymmetrically—to changes in user composition, as indicated by diverging elasticities on the male and female sides. These results underscore the role of digital intermediation in shaping modern family formation. Unlike traditional arrangements, platforms like IBJ provide structured, data-rich environments where frictions are minimized, outcomes are verified, and behavioral dynamics—potentially differing across user groups—are observable, enabling new empirical insights.
@@ 2 Data
I use confidential data from IBJ covering the period 2014–2025. In 2024, IBJ accounted for 3.3% of all new-marriages in Japan, with over 10,000 engagements annually. While this share is smaller than administrative marriage registers, IBJ offers unique visibility into the pre-marriage process. The platform collects verified user data—age, income, education, and marital status—and tracks the full sequence of actions: search, proposals, messages, and dates. All records are time-stamped and linked to engagement outcomes. IBJ users are highly selected into serious long-term partnership search. Entry requires screening and substantial upfront, monthly, and exit fees, effectively excluding casual users. In contrast to dating apps with self-reported profiles and open-ended interactions, IBJ is a closed, consultant-mediated two-sided platform where mutual consent leads to confirmed engagements. This structured setting makes the data well-suited for analyzing matching efficiency and responsiveness over time.
@@
Panels (a) and (b) in Figure 1 display trends in the number of active users and engagements on the IBJ platform from 2014 to 2025. Both female $F$ and male $M$ participation rise steadily, with male user growth outpacing that of females—particularly after 2020—leading to a gradual decline in market tightness, measured as the gender balance ratio $\frac{F}{M}$. Alongside this compositional shift, the number of engagements $E$, realized within the platform, remains relatively flat until around 2020, after which it increases sharply. This acceleration likely reflects improvements in platform performance, such as better algorithmic matching, increased activity, or broader adoption of digital matching services.
Panel (c) illustrates partner finding rates for female and male users per month, measured respectively as $\frac{E}{F}$ and $\frac{E}{M}$. Both rates increase over time, with females consistently achieving higher matching rates. This suggests improved match formation efficiency, particularly for female users, despite the growing supply of male participants. The male rate displays greater month-to-month volatility, indicating more sensitivity to market conditions.
@@ 3 Empirical Framework
Let $E_{t}$ denote the number of within-platform engagements at time $t$, $F_{t}$ and $M_{t}$ the number of active female and male users, and $A_{t}$ a time-varying matching efficiency parameter. Efficiency $A_{t}$ measures how effectively the marriage market turns available men and women into engagements—for a given number of men and women, higher efficiency means more couples are formed. The matching efficiency can be arbitrarily interpreted as search effort, app-based improvements in user interface and experience, and so on (on the female side). Let $(A,F,M)$ denote random variables corresponding with realizations subscripted by time $t$. Let assume that $M$ and $A$ are independent conditional on $F$, that is $M\mathop{\perp}A|F$. I assume a matching function of the form $E_{t}=m(A_{t}F_{t},M_{t})$ with constant returns to scale, which is typically assumed in the literature. Given the assumptions, by applying the nonparametric identification results of Matzkin (2003) to avoid functional form assumptions, Proposition 1 of Lange and Papageorgiou (2020) shows that the observed joint distribution of engagements, female users, and male users, denoted as $G(E,F,M)$, nonparametrically identifies the joint distribution of matching efficiency and female participation, $H(A,F)$, as well as the matching function $m(AF,M):\mathbb{R}^{2}_{+}\rightarrow\mathbb{R}_{+}$, up to a normalization of the efficiency parameter $A$ at a given point $A_{0}$ in the support of $(A,F,M)$.
@@ 4 Results
Using nonparametric methods proposed by Lange and Papageorgiou (2020), I estimate time-varying matching efficiency $A_{t}$ and elasticities with respect to effective female search input $A_{t}F_{t}$ and male users $M_{t}$. [^4] Details are shown in Appendix.
@@ 4.1 Matching efficiency and elasticity in the marriage platform
In Figure 2, matching efficiency increases threefold from baseline levels in 2014, with gradual acceleration beginning around 2017 and continuing through the post-2020 period. Notably, efficiency nearly triples between 2017 and 2025. This improvement may be partly attributable to structural changes within the platform, including the consolidation of other major marriage agencies into the IBJ network. For example, the integration of ZWEI—another nationwide marriage consultation service—into the IBJ Group in 2020 likely expanded the effective user base and improved cross-platform coordination. [^5] Source: Nikkei (2020), “IBJ to Acquire Zwei to Expand Marriage Services,” March 11. Such developments may have enhanced matching opportunities and the sophistication of recommendation mechanisms, contributing to substantial gains in match formation efficiency.
Next, we calculate matching elasticity, which captures how sensitive the number of new couples is to changes in the number of available men or women. Elasticity is estimated by a lasso regression projecting matches on fully interacted 2nd order polynomials of $AF$ and $M$. Elasticity with respect to female-side input ranges from 0.5 to 0.9, while male-side elasticity ranges from 0.5 to 0.3, reflecting an increasingly asymmetric impact of each side on overall matching outcomes.
Together, these trends indicate that the IBJ platform has become not only more efficient in producing engagements, but also more responsive—albeit asymmetrically—to changes in the user base composition. The rise in efficiency and divergence in elasticities suggest improved functionality in the underlying matching process, potentially driven by better recommendation algorithms, a more diverse user base, or targeted engagement strategies differentiated by gender.
@@
4.2 Regional differences in matching efficiency and elasticity
@@ 5 Conclusion
This paper documents substantial improvements in matching efficiency on a large-scale marriage platform in Japan. Leveraging verified user data and well-defined engagement outcomes, I apply a nonparametric approach to estimate time-varying matching functions. The estimates show that the platform has become more effective and increasingly asymmetric in its responsiveness to user-side dynamics, with greater sensitivity to changes on the female side. Regional analysis confirms consistent functionality across major urban areas. These findings highlight the potential of digital matching platforms to reshape traditional partner search markets through structured intermediation and algorithmic improvements tailored to different sides of the market.
@@ A.1 Estimation details
where $\kappa(\cdot)$ is a bivariate normal kernel function with bandwidth set to 0.75.
@@
Finally, I compute local elasticities of the matching function by 2nd order LASSO regression of engagements $E$ on female and male user counts, interacted with the estimated effective search input $AF$. This allows us to estimate the marginal responsiveness of engagements to changes in female and male participation:
$\text{Elasticity w.r.t. }F:\quad\frac{d\log m(AF,M)}{d\log F}=\frac{d\log m(AF,M)}{d\log AF},$